OpenStudy (anonymous):

Solve by completing the square. x^2+12x-8=0

OpenStudy (cwrw238):

you first transform the equation to the form (x + a)^2 + b = 0

OpenStudy (wolf1728):

To complete the square: Move the "non X" term to the right: x^2+12x-8=0 x² + 12x = 8 Divide the equation by the coefficient of X² which in this case is 1 so there's no change x² + 12x = 8 Now for "completing the square" we: take the coefficient of X which is 12 divide that by 2 6 square that number 36 then add it to both sides of the equation x² + 12x +36 = 8 +36 Take the square root of BOTH sides of the equation (x+6) = sqrt(44) x = sqrt(38) x = 6.164414003 x = -6.164414003

OpenStudy (anonymous):

Thanks bro

OpenStudy (wolf1728):

u r welcome